# End-to-End Probabilistic Framework for Learning with Hard Constraints

### v1

@@ 3.4 DPPL for Various Constraint Types
In this subsection, we discuss how to compute the posterior mean and covariance in Equation 8 arising from the DPPL for various constraint types (linear equality, nonlinear equality, and convex inequality), where either a closed-form expression exists or optimization methods are needed to approximate it. See Table 5 in Appendix C for a summary.
@@ 3.4.1 Linear Equality Constraints
For linear equality constraints, $h(\hat{u})=A\hat{u}-b=0$, where $A\in\mathbb{R}^{q\times n},q\leq n$, has full row rank $q$, we can derive a closed-form solution to the constrained least squares Problem 7. This is given by the following proposition, the proof of which can be found in Section C.1.
Proposition 3.1.
For linear equality constraints, $h(\hat{u})=A\hat{u}-b=0$, with $A\in\mathbb{R}^{q\times n}$, with full row rank $q$, where $q\leq n$, and $b\in\mathbb{R}^{q}$, the optimal solution $u^{*}$ to Problem 7 is given as $u^{*}(z)=P_{Q^{-1}}z+(I-P_{Q^{-1}})A^{{\dagger}}b$, where $P_{Q^{-1}}=I-Q^{-1}A^{\top}(AQ^{-1}A^{\top})^{-1}A,$ denotes an oblique projection operator, and $A^{{\dagger}}$ denotes the Moore-Penrose inverse. In addition, if $\mathbf{Z}\sim\mathcal{F}(\mu,\Sigma)$ and $z\sim\mathbf{Z}$ for multivariate, location-scale distribution $\mathcal{F}$, then $u^{*}\sim\mathbf{Y}$, where $\mathbf{Y}\sim\mathcal{F}(\hat{\mu},\hat{\Sigma})$ and $\hat{\mu},\hat{\Sigma}$ are given in Equation 8 with $\mathcal{T}(z)=u^{*}(z)$, which simplifies to the closed-form expressions, $\hat{\mu}=P_{Q^{-1}}\mu+(I-P_{Q^{-1}})A^{{\dagger}}b$ and $\hat{\Sigma}=P_{Q^{-1}}\Sigma P^{\top}_{Q^{-1}}$.
Linear equality constraints occur in a wide range of applications, including coherency constraints in hierarchical time series forecasting (Hyndman et al., 2011; Rangapuram et al., 2021; Petropoulos et al., 2022; Olivares et al., 2024b), divergence-free conditions in incompressible fluid flows (Raissi et al., 2019; Richter-Powell et al., 2022), boundary conditions in PDEs (Saad et al., 2023), and global linear conservation law constraints (Hansen et al., 2023; Mouli et al., 2024).
@@ 3.4.2 Nonlinear Equality Constraints
For nonlinear equality constraints, $h(u)=0$, we can no longer derive the exact closed-form expression for the solution. Instead, we can provide an expression which is satisfied by the optimal solution. This can then be solved for the posterior mean $\hat{\mu}=u^{*}(\mu)$ in Equation 8 with the nonlinear transformation $\mathcal{T}(\mu)=u^{*}(\mu)$ with iterative optimization methods, e.g., Newton’s Method. (We can then compute the posterior covariance $\hat{\Sigma}$ in Equation 8 by estimating the Jacobian $J_{\mathcal{T}}(\mu)$ by differentiating the nonlinear equations $u^{*}(z)=z-Q^{-1}\nabla h(u^{*}(z))^{\top}\lambda\,,\ h(u^{*}(z))=0$ with respect to $z$ via the implicit function theorem (Blondel et al., 2022), and evaluating it at $\mu$. See Gould et al. (2022) for an extensive review to compute the derivatives of optimization programs.) The equations satisfied by the exact solution are given in the following proposition, the proof of which is given in Section C.2.
Proposition 3.2.
For nonlinear equality constraints, $h(\hat{u})=0\in\mathbb{R}^{q}$, where $h:\mathbb{R}^{n}\to\mathbb{R}^{q}$, the optimal solution $u^{*}(z)$ to Problem 7 forms a pair $(u^{*}(z),\lambda^{*})$ which satisfies $u^{*}(z)=z-Q^{-1}\nabla h(u^{*}(z))^{\top}\lambda^{*}$ and $h(u^{*}(z))=0$. In addition, if $\mathbf{Z}\sim\mathcal{F}(\mu,\Sigma)$ and $z\sim\mathbf{Z}$ for multivariate, location-scale distribution $\mathcal{F}$, then $u^{*}\sim\mathbf{Y}$, where $\mathbf{Y}\sim\mathcal{F}(\hat{\mu},\hat{\Sigma})$ and $\hat{\mu},\hat{\Sigma}$ are given in Equation 8 with $\mathcal{T}(z)=u^{*}(z)$.
Nonlinear equality constraints naturally arise in applications that involve structural, physical, or geometric consistency. These include closed-loop kinematics in robotics (Toussaint et al., 2019), nonlinear conservation laws (LeVeque, 1990) in PDE-constrained surrogate modeling (Biegler et al., 2003; Zahr and Persson, 2016; Négiar et al., 2023) with applications in climate modeling (Bolton and Zanna, 2019; Zanna and Bolton, 2020; Beucler et al., 2021), compressible flows in aerodynamics (Tezaur et al., 2017) and atomic modeling (Müller, 2022; Sturm and Wexler, 2022).
@@
For convex inequality constraints, $\hat{u}$ in Problem 7 is in a convex set, $\mathcal{C}\subset\mathbb{R}^{n}$. Closed-form expressions (such as those in previous subsections for linear and nonlinear equality constraints) do not exist (Boyd and Vandenberghe, 2004). Instead, we rely on convex optimization solvers to ensure computational efficiency and scalability to compute the solution $u^{*}$. The gradients of the convex program can be calculated efficiently using sensitivity analysis (Bertsekas, 1997; Bonnans and Shapiro, 2013), argmin differentiation (Sun et al., 2022; Agrawal et al., 2019; Amos and Kolter, 2017; Gould et al., 2016), and/or variational analysis (Rockafellar and Wets, 2009). These techniques provide a means to compute the Jacobian $J_{\mathcal{T}}(\mu)$, which represents the sensitivity of the optimal solution $u^{*}$ to changes in the input vector $\mu$, whose projection we are essentially computing to the convex constraints space. See Section C.3 for details.
Convex inequality constraints arise naturally in many scientific and engineering applications. For example, total variation (TV) regularization is widely used to promote smoothness or piecewise-constant structure in spatial fields, e.g., image denoising (Rudin et al., 1992; Boyd and Vandenberghe, 2004) and total variation diminishing (TVD) constraints to avoid spurious artificial oscillations in numerical solutions to PDEs (Harten, 1997; Tezaur et al., 2017; Schein et al., 2021). Other common convex constraints include box constraints, which enforce boundedness of physical or operational quantities (Bertsekas, 1997).
@@ 3.5 Sample-free with Closed-form CRPS
[…]
@@ 4 Empirical Results
    Does training with the distributiuon-agnostic proper scoring rule, CRPS instead of NLL, improve performance?
@@
We compare two variants of ProbHardE2E, i.e., ProbHardE2E-Ob, which uses a general oblique projection ($Q=\Sigma^{-1}$) projection and is our default unless otherwise specified, and ProbHardE2E-Or, which uses an orthogonal projection ($Q=I$), against several probabilistic deep learning baselines commonly used for uncertainty quantification in constrained PDEs and probabilistic time series forecasting. For PDEs, ProbHardE2E uses VarianceNO (Mouli et al., 2024) as the probabilistic base model that extends the deterministic SciML Fourier Neural Operator (FNO) (Li et al., 2021) model to handle probabilistic estimates by augmenting the final layer to predict both the mean and variance of the solution (Lakshminarayanan et al., 2017). We compare ProbHardE2E with: (i) HardC, which is based on Négiar et al. (2023); Hansen et al. (2023), and which imposes the orthogonal projection only on the mean, but does not update the covariance; (ii) ProbConserv (Hansen et al., 2023; Mouli et al., 2024), which applies the oblique projection only at inference time, and works only with linear constraints (in the nonlinear constraint case, we compare with ProbHardInf, which is a variant of ProbConserv that imposes the nonlinear constraint at inference time only); (iii) SoftC (Hansen et al., 2023), which introduces a soft penalty on constraint violation à la PINNs (Raissi et al., 2019; Li et al., 2024) during training but does not guarantee constraint satisfaction at inference; and (iv) the unconstrained model backbone VarianceNO. For hierarchical time-series forecasting, ProbHardE2E uses DeepVAR (Salinas et al., 2019) as the probabilistic base model. We compare ProbHardE2E with: (i) ProbConserv; (ii) HierE2E (Rangapuram et al., 2021), which enforces linear constraints via an end-to-end orthogonal projection; classical statistical approaches including ARIMA-NaiveBU, ETS-NaiveBU  (Hyndman et al., 2011), (iii) PERMBU-MINT (Taieb et al., 2017); and (iv) the unconstrained model backbone DeepVAR.
@@
We evaluate ProbHardE2E using the following metrics: Mean Squared Error (MSE) for measuring the mean prediction accuracy; Constraint Error (CE) (conservation law for PDEs and coherency for hierarchical time series forecasting), which measures the error in the constraints; and Continuous Ranked Probability Score (CRPS), which is the most meaningful metric that measures performance in uncertainty quantification (UQ). (See Section H.3 for details on the metrics.) For each model, we report these metrics when trained with either CRPS or Negative Log-Likelihood (NLL) as the loss. Although originally optimized with NLL, we also train a CRPS-based variant of ProbConserv to ensure a fair comparison. The experiments are conducted on a single NVIDIA V100 GPU in the PDEs case, and on an Intel(R) Xeon(R) CPU E5-2603 v4 @ 1.70GHz in the time series forecasting case. To ensure scalability, we use a diagonal covariance matrix $Q$ in Problem 7, following prior work (Hansen et al., 2023; Mouli et al., 2024). (See Appendix E for alternative structures, e.g., low-rank and full covariances.)
@@ 4.1 Linear Conservation and Hierarchical Constraints
Table 1: Test metrics on constrained PDEs across four datasets, which are ordered top to bottom in their learning difficulty. Metrics include MSE ×$\times 10^{-5}$, constraint (conservation) error (CE) ×$\times 10^{-3}$, and CRPS ×$\times 10^{-3}$. Each algorithm is trained with either CRPS or NLL. Best values per row are highlighted in bold.
@@
Table 2: CRPS ×$\times 10^{-3}$ for hierarchical time‑series datasets across various probabilistic forecasting algorithms. Constraint (coherency) error (CE) is given in parathesis and is equal to 0 for all methods except the base unconstrained DeepVAR. PERMBU-MINT is not available for Tourism-L, because the dataset has a nested hierarchical structure, and PERMBU-MINT is not well-defined on this type of dataset (Rangapuram et al., 2021). ARIMA-NaiveBU, ETS-NaiveBU and PERMBU-MINT are deterministic models with no model uncertainty.
@@
Labour 36.1$\pm$ 2.7 (0) 28.6$\pm$6.5 (0) 45.8$\pm$6.5 (0) 50.5$\pm$20.6 (0) 45.3 (0) 43.2 (0) 39.3 (0) 38.2$\pm$4.5 (0.215)
Tourism 98.9$\pm$13.0 (0) 82.4$\pm$6.6 (0) 100.7$\pm$7.7 (0) 103.1$\pm$16.3 (0) 113.8 (0) 100.8 (0) 77.1 (0) 92.5$\pm$2.2 (2818.01)
Tourism-L 155.2$\pm$3.6 (0) 156.4$\pm$9.4 (0) 176.9$\pm$21.5 (0) 161.3$\pm$10.9 (0) 174.1 (0) 169.0 (0) 158.1$\pm$10.2 ($7\times 10^{4}$)
Traffic 55.0$\pm$10.6 (0) 60.6$\pm$7.8 (0) 71.0$\pm$3.9 (0) 41.8$\pm$7.8 (0) 80.8 (0) 66.5 (0) 67.7 (0) 40.0$\pm$2.6 (0.192)
Wiki 212.1$\pm$29.4 (0) 215.8$\pm$16.9 (0) 264.7$\pm$ 30.7 (0) 216.5$\pm$26.7 (0) 377.2 (0) 467.3 (0) 281.2 (0) 229.4$\pm$15.8 (8398.55)
@@
Regarding (Q3) on the training objective, Table 1 shows that training with the proper scoring rule CRPS as the loss function improves UQ (measured by CRPS) across nearly all PDE datasets and SciML models, compared to training with the commonly-used NLL in previous SciML works. The only exception is HardC in the PME. The CRPS training objective also improves mean accuracy (measured by MSE) in approximately three-quarters of the dataset-model experiments. In addition, Table 2 shows that on four out of five time series datasets, we are able to improve upon the results of Hier-E2E, which uses the DeepVAR base model with an orthogonal projection on the samples, and which optimizes the sampling-based quantile loss by projecting directly on the distribution parameters and using the closed-form CRPS.
@@
[…]
4.2 Extension to General Nonlinear Equality and Convex Inequality Constraints (Q5)
In this subsection, we test ProbHardE2E on more challenging constraints types, i.e., nonlinear equality and convex inequality, respectively, to address question (Q5) on PDE datasets, as (current) time series forecasting applications usually need predictions to satisfy only linear (e.g., hierarchical) constraints.
@@ 4.2.1 Nonlinear Equality Constraints
We test ProbHardE2E with general nonlinear constraints using nonlinear conservation laws from PDEs. (See Section I.2 for details.) Table 3 shows the superior performance of enforcing nonlinear constraints with ProbHardE2E. We see an even larger MSE performance improvement of 1517$\approx$ 15-17$\times$ 15 - 17 × when trained on CRPS, and CRPS performance improvement of 2.5$\approx$ 2.5$\times$ 2.5 × over the various baselines that apply the nonlinear constraint just at inference time (ProbHardInf), as a reduced linear constraint (ProbConserv) at inference time, and the unconstrained model (VarianceNO). These results especially highlight the benefit of training end-to-end with the constraint in the nonlinear case. In addition, LABEL:fig:nonlin_pme shows that ProbHardE2E is able to significantly better capture the shock and has tighter uncertainty estimates, leading to lower CRPS values than the baselines.
Table 3: Test metrics on the nonlinear PME with PDE coefficient $k(u)=u^{m}$, which controls the sharpness of the solution (larger values are “harder”), for NLL and CRPS training. The training and test parameters are sampled from this range of $m$. Metrics include MSE ×$\times 10^{-6}$, calibration error (CE) ×$\times 10^{-3}$, and CRPS ×$\times 10^{-4}$. Best values per row and metric are bolded.

### v2

@@ 3.4 DPPL for Various Constraint Types
In this subsection, we discuss how to compute the DPPL for various constraint types (linear equality, nonlinear equality, and convex inequality) for both train and inference modes. Table 1 shows these constraints types require different treatments: linear equality have closed-form projections, nonlinear equality can be solved with iterative methods, and convex inequality require optimization solvers.
Table 1: Summary of DPPL in ProbHardE2E for various constraint types. For linear equality constraints, the oblique projection $P_{Q^{-1}}=I-Q^{-1}A^{\top}(AQ^{-1}A^{\top})^{-1}A$; for nonlinear equality constraints, $R$ denotes the first-order optimality conditions.
Constraint Type Solution $u^{*}(z)$ Solver Type Jacobian $J_{\mathcal{T}}$
Linear Equality $P_{Q^{-1}}z+(I-P_{Q^{-1}})A^{{\dagger}}b$ closed-form $P_{Q^{-1}}$
Nonlinear Equality $(u^{*},\lambda^{*})$ s.t. $R(u^{*},\lambda^{*};z)=0$ nonlinear implicit differentiation
sensitivity analysis;
argmin differentiation
@@ 3.4.1 Linear Equality Constraints
For linear equality constraints, $h(\hat{u})=A\hat{u}-b=0$, where $A\in\mathbb{R}^{q\times n},q\leq n$, has full row rank $q$, we can derive a closed-form solution to the constrained least squares Problem 7. In this case, both training and inference modes are equivalent since the DPPL projection is exact. (See Section C.1.)
@@ 3.4.2 Nonlinear Equality Constraints
For nonlinear equality constraints, $h(\hat{u})=0$, we can no longer derive the exact closed-form expression for the solution. Instead, we can provide an expression which is satisfied by the optimal solution. In particular, we approximate the parameter-level projection at training time. This can then be solved for the posterior mean $\hat{\mu}=u^{*}(\mu)$ in Equation 8 with the nonlinear transformation $\mathcal{T}(\mu)=u^{*}(\mu)$ with iterative optimization methods, e.g., Newton’s Method. (We can then compute the posterior covariance $\hat{\Sigma}$ in Equation 8 by estimating the Jacobian $J_{\mathcal{T}}(\mu)$ by differentiating the nonlinear equations $u^{*}(z)=z-Q^{-1}\nabla h(u^{*}(z))^{\top}\lambda\,,\ h(u^{*}(z))=0$ with respect to $z$ via the implicit function theorem (Blondel et al., 2022), and evaluating it at $\mu$. (See Section C.2.) At inference, we project each sample exactly with our custom, batched optimization solver to ensure strict constraint feasibility.
@@
For convex inequality constraints, $\hat{u}$ in Problem 7 is in a convex set, $\mathcal{C}\subset\mathbb{R}^{n}$. Closed-form expressions (such as those in previous subsections for linear and nonlinear equality constraints) do not exist (Boyd and Vandenberghe, 2004). Instead, we rely on convex optimization solvers to ensure computational efficiency and scalability to compute the solution $u^{*}$ in training. The gradients of the convex program can be calculated efficiently using sensitivity analysis (Bertsekas, 1997; Bonnans and Shapiro, 2013), argmin differentiation (Sun et al., 2022; Agrawal et al., 2019; Amos and Kolter, 2017; Gould et al., 2016), and/or variational analysis (Rockafellar and Wets, 2009). These techniques provide a means to compute the Jacobian $J_{\mathcal{T}}(\mu)$, which represents the sensitivity of the optimal solution $u^{*}$ to changes in the input vector $\mu$, whose projection we are essentially computing to the convex constraints space. During inference, we solve the convex program per sample. (See Section C.3.)
@@ 3.5 Sample-free with Closed-form CRPS
We use a closed-form expression for the CRPS to enable a computationally efficient and sample-free approach for evaluating the CRPS in the loss function $\ell$ in Equation 3. Calculating the CRPS for an arbitrary distribution requires generating samples (Rangapuram et al., 2021; Gneiting and Raftery, 2007), but closed-form expressions for the CRPS exist for several location-scale distributions (Gaussian, logistic, student’s t, beta, gamma, uniform). Most notably, for the univariate Gaussian, the closed-form CRPS is given as: $\text{CRPS}_{\mathcal{N}}(z)=\left[z\cdot\left(2\mathrm{P}(z)-1\right)+2p(z)-\frac{1}{\sqrt{\pi}}\right],$ where $p(z)=\frac{1}{\sqrt{2\pi}}\exp{(-z^{2}/2)}$ denotes the standard normal probability density function (PDF), and $\mathrm{P}(z)=\int_{-\infty}^{z}p(y)dy$ denotes the standard normal cumulative distribution function (CDF) for $z\sim\mathcal{N}(0,1)$ (Gneiting et al., 2005; Taillardat et al., 2016). This sample-free formulation is especially beneficial when the DPPL is computationally intensive, e.g., in the presence of nonlinear constraints.
@@ 4 Empirical Results
    Does training with the distribution-agnostic proper scoring rule, CRPS instead of NLL, improve performance?
@@
We compare two variants of ProbHardE2E, i.e., ProbHardE2E-Ob, which uses a general oblique projection ($Q=\Sigma^{-1}$) projection and is our default unless otherwise specified, and ProbHardE2E-Or, which uses an orthogonal projection ($Q=I$), against several probabilistic deep learning baselines commonly used for uncertainty quantification in constrained PDEs and probabilistic time series forecasting. For PDEs, ProbHardE2E uses VarianceNO (Mouli et al., 2024), which is a probabilisitic extension of the Fourier Neural Operator (FNO) (Li et al., 2021) as the unconstrained model. We compare ProbHardE2E with: (i) HardC, which is based on Négiar et al. (2023); Hansen et al. (2023), and which imposes the orthogonal projection only on the mean, but does not update the covariance; (ii) ProbConserv (Hansen et al., 2023), which applies the oblique projection only at inference time, and works only with linear constraints (in the nonlinear constraint case, we compare with ProbHardInf, which is a variant of ProbConserv that imposes the nonlinear constraint at inference time only); (iii) SoftC (Hansen et al., 2023), which introduces a soft penalty on constraint violation à la PINNs (Raissi et al., 2019; Li et al., 2024) during training but does not guarantee constraint satisfaction at inference; and (iv) the unconstrained model backbone VarianceNO. For hierarchical time-series forecasting, ProbHardE2E uses DeepVAR (Salinas et al., 2019) as the probabilistic base model. We compare ProbHardE2E with: (i) ProbConserv; (ii) HierE2E (Rangapuram et al., 2021), which enforces linear constraints via an end-to-end orthogonal projection; classical statistical approaches including ARIMA-NaiveBU, ETS-NaiveBU  (Hyndman et al., 2011, 2025), (iii) PERMBU-MINT (Taieb et al., 2017; Olivares et al., 2022); and (iv) the unconstrained model backbone DeepVAR.
@@
We evaluate ProbHardE2E using the following metrics: Mean Squared Error (MSE), which measures the mean prediction accuracy; Constraint Error (CE), which measures the constraint errors on the samples (conservation law for PDEs and coherency for hierarchical time series forecasting); and Continuous Ranked Probability Score (CRPS), which measures performance in uncertainty quantification (UQ). (See Section H.3 for details on the metrics.) For each model, we report these metrics when trained with either CRPS or Negative Log-Likelihood (NLL) as the loss. Although originally optimized with NLL, we also train a CRPS-based variant of ProbConserv to ensure a fair comparison. The experiments are conducted on a single NVIDIA V100 GPU in the PDEs case, and on an Intel(R) Xeon(R) CPU E5-2603 v4 @ 1.70GHz in the time series forecasting case. To ensure scalability, we use a diagonal covariance matrix $Q$ in Problem 7, following prior work (Hansen et al., 2023; Mouli et al., 2024). (See Appendix E for low-rank and full covariance structures.)
@@ 4.1 Linear Conservation and Hierarchical Constraints
Table 2: Test metrics on constrained PDEs across four datasets, which are ordered top to bottom in their learning difficulty. Metrics include MSE $\times 10^{-5}$, constraint (conservation) error (CE) $\times 10^{-3}$, and CRPS $\times 10^{-3}$. Each algorithm is trained with either CRPS or NLL. Best values per row are highlighted in bold.
@@
Table 3: CRPS $\times 10^{-3}$ for hierarchical time‑series datasets across various probabilistic forecasting algorithms. Constraint (coherency) error (CE) is given in parenthesis and is equal to 0 for all methods except the base unconstrained DeepVAR. PERMBU-MINT is not available for Tourism-L, because the dataset has a nested hierarchical structure, and PERMBU-MINT is not well-defined on this type of dataset (Rangapuram et al., 2021). ARIMA-NaiveBU, ETS-NaiveBU and PERMBU-MINT are deterministic models with no model uncertainty.
@@
[…]
@@
Regarding (Q3) on the training objective, Table 2 shows that training with the proper scoring rule, such as CRPS, improves UQ (measured by CRPS) across nearly all PDE datasets compared to training with the commonly-used NLL in previous SciML works. The only exception is HardC in the PME. The CRPS training objective also improves mean accuracy (measured by MSE) in approximately three-quarters of the dataset-model experiments. In addition, Table 3 shows that on four out of five time series datasets, we improve upon the results of Hier-E2E, which uses the DeepVAR base model with an orthogonal projection on the samples, and which optimizes the sampling-based quantile loss by projecting directly on the distribution parameters.
@@
Regarding (Q4) on the computational efficiency of our sampling-free approach, LABEL:fig:pde_timing shows the training time per epoch for the hierarchical time series datasets. Models trained for time series and PDEs (see Section I.1) with 100 posterior samples per training step incur a $3.3\text{--}4.6\times$ increase in epoch time relative to ProbHardE2E, which avoids sampling altogether by using a closed-form CRPS loss. Note that the computational overhead of ProbHardE2E is approximately $2\times$ that of the unconstrained model. However, compared to the sampling-based probabilistic baselines, our approach with the DPPL layer that directly projects distribution parameters and a closed-form CRPS objective offers significant training-time speed-ups across all forecasting and PDE testbeds.
4.2 General Nonlinear Equality and Convex Inequality Constraints (Q5)
In this subsection, we test ProbHardE2E on nonlinear equality and convex inequality constraints to address question (Q5) on PDE datasets, as (current) time series forecasting applications usually need predictions to satisfy only linear (e.g., hierarchical) constraints.
@@ 4.2.1 Nonlinear Equality Constraints
We test ProbHardE2E with general nonlinear constraints using nonlinear conservation laws from PDEs. (See Section I.2 for details.) Importantly, Table 4 shows that even in this challenging case of nonlinear constraints, the constraint error (CE) on the samples is 0 so that we ensure strict feasibility on the samples. In addition, we see superior performance of enforcing nonlinear constraints with ProbHardE2E. We see an even larger MSE performance improvement of $\approx 15-17\times$ when trained on CRPS, and CRPS performance improvement of $\approx 2.5\times$ over the various baselines that apply the nonlinear constraint just at inference time (ProbHardInf), as a reduced linear constraint (ProbConserv) at inference time, and the unconstrained model (VarianceNO). These results highlight the benefit of training end-to-end with the constraint in the nonlinear case. In particular, this validates our dual mode training and inference approach, where we obtain the computational benefits of projecting on the parameters and approximating satisfying the constraint during training, and exact constraint sanctification at inference time while achieving state-of-the-art performance measured in CRPS. In addition, LABEL:fig:nonlin_pme shows that ProbHardE2E is able to significantly better capture the shock and has tighter uncertainty estimates, leading to lower CRPS values than the baselines.
(a) ProbHardE2E on the various constraint types. (a) Linear Equality: Average time per iteration (in seconds) for ProbHardE2E, compared to the HierE2E on five hierarchical time-series datasets; (b) Nonlinear Equality: Mean $\pm$3 standard deviation for the PME with conservation constraint at time $t=0.51$, with PDE parameter $m_{\text{train}}\in[3,4]$ and $m_{\text{test}}=3.88$; (c) Convex Inequality: Mean $\pm$3 standard deviation for linear advection with TVD constraint at time $t=0.51$, with PDE parameter $\beta_{\text{train}}\in[1,2]$ and $\beta_{\text{test}}=1.5$. The horizontal axes in (b)-(c) are zoomed in to highlight the uncertainty near the propagating front.
Table 4: Test metrics on the nonlinear PME with PDE coefficient $k(u)=u^{m}$, which controls the sharpness of the solution (larger values are “harder”), for NLL and CRPS training. The training and test parameters are sampled from this range of $m$. Metrics include MSE $\times 10^{-6}$, calibration error (CE) $\times 10^{-3}$, and CRPS $\times 10^{-4}$. Best values per row and metric are bolded.